I am not sure if I picked the right catagory but I am having a hard time with on algebra word problem.
If the formula R = -0.037t + 50.1 can be used to predict the world record in the 400 - meter dash t years after 1925, for what years will the world records be 47.6 seconds or less?
If you could help me with this I would SO appreciate it!
Jen

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

You are given:
R = -0.037t + 50.1
where
R is the record
t is years after 1925
.
So, they want to know when the record will be 47.6 seconds or less...
So, just "plug it in" and solve for 't':
R = -0.037t + 50.1
47.6 = -0.037t + 50.1
47.6 + 0.037t = 50.1
0.037t = 2.5
t = 2.5/0.037
t = 67.568 years after 1925
.
So,
1925 + 67.568 = 1992.568
.
Conclusion:
1992 and AFTER the world record will be 47.6 seconds or less



Answer by stanbon(75887)

If the formula R = -0.037t + 50.1 can be used to predict the world record in the 400 - meter dash t years after 1925, for what years will the world records be 47.6 seconds or less?
----------------------------------------
EQUATION:
47.6 >= -0.037t + 50.1
-0.037t <= 47.6-50.1
-0.037t <= -2.5
t >= 67.567
--------------
Year?
year >= 1925+67.567 >= 1993
=====================
Cheers,
Stan H.